# Difficult control is related to instability in biologically inspired Boolean networks

### v1

@@ 4 Discussion
With a few notable exceptions, our previous work indicated that the average required level of control primarily depends on the number of preexisting attractors and shows no dependence on the system’s size. While these exceptions were not numerous enough to undermine the general trend of ‘easy controllability’, they raised the question of whether biological systems exist that are inherently harder to control.
This work represents a more in-depth analysis of these exceptions. Similarly to Ref. [4], we compare the results from our database of biological networks to those obtained in several ensembles of systems with randomly assigned network topologies and dynamic rules. These random ensembles are particularly important to include in that they produce a larger number of harder to control networks. Additionally, we roughly doubled the number of biological networks we tested (increasing from 49 to 104 models) through the inclusion of networks cataloged in Ref. [14]. Importantly, no new exceptions to our original scaling law were observed.
Our primary observation here unveils a consistent pattern in these exceptions, found in both the biological and random ensembles. This pattern originates from the presence of isolated fixed points—fixed points that attract no other states and prove to be among the attractors most challenging to control, with a control kernel that often includes all or almost all core nodes in the network.
@@
Unfortunately, this connection between unstable fixed points and difficult control itself has exceptions. In particular, it is possible to construct networks that are difficult to control without having globally unstable fixed points. [^2] We describe one such exception in the SI, a network dynamic in which a locally unstable fixed point has maximally distant network states that map back to it. We find only one similar network in the random ensembles that we study, and we conjecture that such contrived examples are unlikely in real biological systems. Still, by efficiently counting the unstable fixed points of a network, we are able to better explain the variance in control kernel sizes by incorporating the prevalence of these unstable fixed points. This constitutes a useful tool to predict the expected amount of required control over a network based solely on partial information of the attractor landscape, thus avoiding the NP-hard problem of evaluating individual control kernels.
@@
The observation that difficult control is related to unstable fixed points is reassuring for several reasons. Firstly, these fixed points likely result from the strict deterministic nature of updates in the considered biological models. Most sources of stochasticity [27] would likely push the dynamics onto a path leading to a more stable attractor characterized by an adjacent basin of non-negligible size. More importantly, this same property makes it highly improbable for such states to carry biological significance, i.e., to represent actual cell types. Their persistence would require fine-tuned preservation of an unstable dynamical state.
@@
The biological irrelevance of such unstable fixed points is also significant for another reason. As they reintroduce an explicit dependence on the network size, these exceptions would contradict our main statement that control does not scale with the size of the system [4], in direct disagreement with recent results in the genetic reprogramming of mammalian cells [13]. In this light, we interpret our correction (Equation 2) as an empirical method for obtaining a more precise estimate of the required control when employing a simplistic, deterministic model of the regulatory dynamics. Verifying that our size-independent scaling law remains valid in models of genetic regulation that include stochasticity is needed but beyond the scope of this work.
@@
Constructed exceptional network
@@
name size core nodes mean CK size attr. isol. f.p.
$n$ $n_{\mathrm{c}}$ $|CK|\right>$ $r$ $s$
Arabidopsis Thaliana Cell Cycl… 14 14.0 0.0 1 0
B Cell Differentiation 22 20.0 5.9 61 0
BT474 Breast Cell Line Long-te… 25 25.0 14.0 3344 256
BT474 Breast Cell Line Short-t… 16 16.0 9.5 253 24
Body Segmentation In Drosophil… 17 12.0 3.4 10 0
Budding Yeast Cell Cycle 20 20.0 4.8 33 0
Budding Yeast Cell Cycle 2009 18 18.0 0.0 1 0
CD4+ T Cell Differentiation An… 18 18.0 8.6 324 0
Cardiac Development 15 15.0 2.8 6 0
Cell Cycle Transcription By Co… 9 9.0 1.0 2 0
Cortical Area Development 5 5.0 1.0 2 0
Death Receptor Signaling 28 25.0 4.7 97 0
FGF Pathway Of Drosophila Sign… 23 9.0 9.0 512 512
Fanconi Anemia And Checkpoint … 15 15.0 0.0 1 0
Guard Cell Abscisic Acid Signa… 44 37.0 5.4 95 0
HCC1954 Breast Cell Line Long-… 25 25.0 17.8 9452 1504
HCC1954 Breast Cell Line Short… 16 16.0 9.0 274 10
HH Pathway Of Drosophila Signa… 24 13.0 13.0 8192 8192
Human Gonadal Sex Determinatio… 19 18.0 2.3 3 0
Iron Acquisition And Oxidative… 22 17.0 2.0 4 0
Lac Operon 13 13.0 3.2 9 0
Mammalian Cell Cycle 20 18.0 1.7 3 0
Mammalian Cell Cycle 2006 10 10.0 0.0 1 0
Metabolic Interactions In The … 12 11.0 6.2 40 0
Neurotransmitter Signaling Pat… 16 14.0 2.0 4 0
Oxidative Stress Pathway 19 17.0 1.0 2 0
Predicting Variabilities In Ca… 15 15.0 2.8 6 0
Pro-inflammatory Tumor Microen… 26 26.0 2.5 6 0
Processing Of Spz Network From… 24 6.0 6.0 64 64
Regulation Of The L-arabinose … 13 10.0 4.1 21 0
SKBR3 Breast Cell Line Long-te… 25 25.0 17.3 3185 960
SKBR3 Breast Cell Line Short-t… 16 16.0 10.7 702 104
Septation Initiation Network 31 29.0 9.4 2240 0
T Cell Differentiation 23 23.0 5.0 33 0
T-Cell Signaling 2006 40 10.0 3.1 10 0
T-LGL Survival Network 2011 Re… 18 17.0 2.3 3 0
TOL Regulatory Network 24 10.0 10.0 1024 1024
Toll Pathway Of Drosophila Sig… 11 2.0 2.0 4 4
Treatment Of Castration-Resist… 42 14.0 14.0 16384 16384
Trichostrongylus Retortaeformi… 26 25.0 3.2 12 0
Tumour Cell Invasion And Migra… 32 22.0 3.3 15 0
Vegf Pathway Of Drosophila Sig… 18 8.0 8.0 256 256
Wg Pathway Of Drosophila Signa… 26 23.0 14.0 16384 0
Yeast Apoptosis 73 30.0 13.3 11520 0
@@
PubMed ID size core nodes mean CK size attr. isol. f.p.
$n$ $n_{\mathrm{c}}$ $|CK|\right>$ $r$ $s$
11082279 11 11 2.0 4 0
1753781 8 8 3.4 11 1
19524598 12 12 2.0 4 0
19622164_TGF_alpha 17 9 2.8 7 0
19622164_TGF_beta1 17 9 2.8 7 0
20169167 7 7 1.0 2 0
20221256_manual_mod 11 11 3.6 22 0
20659480 20 13 1.7 4 0
21450717_model_2_brassinostero… 10 3 3.0 8 8
21450717_model_4_2_BRX_integra… 20 17 3.0 14 0
21450717_model_5_1_BRX_integra… 20 17 2.0 4 0
21450717_model_5_2 20 17 2.0 4 0
21563979 13 13 3.2 9 0
21639591 25 25 2.0 4 0
21853041 11 11 2.7 8 0
22192526 9 8 3.0 7 0
23056457 17 16 3.8 12 0
23169817_high_dna_damage 16 16 0.0 1 0
23169817_low_dna_damage 16 16 0.0 1 0
23469179 10 9 3.1 8 0
23630177_metastatic_melanoma 10 9 2.0 4 0
23658556_model_1 14 14 3.8 10 0
23658556_model_2 16 16 3.2 9 0
23658556_model_3 15 15 3.8 9 0
23658556_model_4 17 17 3.0 10 0
23658556_model_5 15 15 3.1 8 0
23658556_model_6 15 15 3.1 8 0
23658556_model_7 17 17 4.5 14 0
23658556_model_8 16 16 4.1 12 0
23658556_model_9 17 17 3.2 9 0
23764028 5 2 2.0 4 4
24376455 10 9 2.0 4 0
24564942_threshold 8 8 2.6 5 0
25163068 30 27 7.3 241 0
25967891 13 12 4.1 10 0
26244885 30 29 8.4 1104 0
26346668 30 2 2.0 4 4
27613445 25 23 3.6 22 0
28187161 30 27 6.8 220 0
28209158 9 9 1.7 3 0
28426669_ARF10_greater_ARF5 16 15 3.3 14 0
28426669_ARF10_smaller_ARF5 16 15 3.3 14 0
28455685 15 15 2.0 4 0
29186334 29 16 5.2 25 0
29596489 23 18 2.2 5 0
30024932 21 16 2.3 6 0
30104572 5 5 2.0 3 0
30148917 18 14 4.0 16 0
30323768 21 16 3.8 18 0
30530226 23 19 2.0 5 0
30546316 30 21 10.6 1056 0
31949240 24 21 5.4 45 0
32054948 27 24 2.2 6 0

### v2

@@ 4 Discussion
Alternative approaches to modeling the control of biological networks exist, and results can vary significantly depending on the goals set for the control. For example, the ability to attain any possible target state in a typical GRN described by a linear, continuous system requires control over more than 80% of the network [6]. Here we focus instead on the discrete, nonlinear case, and we use a definition of control that aims to force the system to a single one of its original attractor states from any initial condition.
With a few notable outliers, our previous work on this type of control indicated that the average required number of control nodes primarily depends on the number of preexisting attractors and shows no dependence on the system’s size. While these outliers were not numerous enough to undermine the general trend of ‘easy controllability’, they raised the question of whether biological systems exist that are inherently harder to control.
This work represents a more in-depth analysis of these outliers. Similarly to Ref. [21], we compare the results from our database of biological networks to those obtained in several ensembles of systems with randomly assigned network topologies and dynamic rules. These random ensembles are particularly important to include in that they produce a larger number of harder to control networks. Additionally, we roughly doubled the number of biological networks we tested (increasing from 49 to 104 models) through the inclusion of networks cataloged in Ref. [25]. Importantly, no new outliers from our original scaling law were observed.
Our primary observation here unveils a consistent pattern in these outlier networks, found in both the biological and random ensembles. This pattern originates from the presence of isolated fixed points—fixed points that attract no other states and prove to be among the attractors most challenging to control, with a control kernel that often includes all or almost all core nodes in the network.
@@
This type of bias is also seen in the few biological networks that are unusually difficiult to control. Among the available biological networks, we found two outliers at the 3-sigma level (triangles in Fig. 4A), both related to the ErbB network in breast cell lines: “SKBR3 Breast Cell Line Long-term ErbB Network” and “HCC1954 Breast Cell Line Long-term ErbB Network.” These networks, which are part of the same study focused on resistance mechanisms in breast cancer treatments [45], all include isolated fixed points. Upon examining the attractors that are hardest to control (those with the maximum control kernel size), we find they consist of states where most nodes are inactive, except for a few nodes that self-sustain activity. This appears to be due to a highly biased network core where the all-zero state is an isolated fixed point, coupled with peripheral nodes that can independently sustain themselves. This combination increases the number of hard-to-control fixed points and thereby the average control kernel size. It is worth noting that this ability of peripheral nodes to sustain their own activity may be an effect of the algorithm used in the original study to infer the network logic.
@@
Unfortunately, this connection between unstable fixed points and difficult control itself has exceptions. In particular, it is possible to construct networks that are difficult to control without having globally unstable (isolated) fixed points. [^3] We describe one such counterexample in the SI (also illustrated in the last row of Fig. 1), a network dynamic in which a locally unstable fixed point has maximally distant network states that map back to it. We find only one similar network in the random ensembles that we study, and that network also contains a globally unstable fixed point. Unsurprisingly, we do not find any network resembling such a contrived counterexample among the biological systems. Still, by efficiently counting the unstable fixed points of a network, we are able to better explain the variance in control kernel sizes by incorporating the prevalence of these unstable fixed points. This constitutes a useful tool to predict the expected amount of required control over a network based solely on partial information of the attractor landscape. While determining the number of attractors in a Boolean network is itself a computationally hard problem (#P-complete [46]), once this is known, our approach avoids the NP-hard problem of evaluating individual control kernels.
@@
The observation that difficult control is related to unstable fixed points is reassuring for several reasons. Firstly, these fixed points are not a general feature of models with non-deterministic updating rules. Most sources of stochasticity [47] would push the dynamics toward more stable attractors characterized by larger basins. A weak form of stochasticity arises in asynchronous updating. In this case the identity of fixed-point attractors remains unchanged, but basin sizes can change. As a result, it is possible that isolated fixed points would no longer be isolated. A stronger form of stochasticity would include bit-flips. In this case, we expect that isolated fixed points would effectively disappear. More importantly, the incompatibility of unstable fixed points with stochasticity makes it highly improbable for such states to carry biological significance, i.e., to represent actual cell types. Their persistence would require fine-tuned preservation of an unstable dynamical state.
@@
The biological irrelevance of such unstable fixed points is also significant for another reason. As they reintroduce an explicit dependence on the network size, these cases of difficult control would contradict our main statement that control does not scale with the size of the system [21], in direct disagreement with recent results in the genetic reprogramming of mammalian cells [23]. In this light, we interpret our correction (Equation 2) as an empirical method for obtaining a more precise estimate of the required control when employing a simplistic, deterministic model of the regulatory dynamics. Verifying that our size-independent scaling law remains valid in models of genetic regulation that include stochasticity is needed but beyond the scope of this work.
@@
5 Code and data availability
The python code used to recreate the analysis and figures, as well as CSV files containing the relevant data, are available as a Zenodo repository under the accession code 10.5281/zenodo.13819680 [https://doi.org/10.5281/zenodo.13819680]. This code depends on a number of open-source software packages: neet, numpy, pandas, and networkx.
6 Acknowledgments
The authors acknowledge Research Computing at Arizona State University for providing HPC resources that have contributed to the research results reported in this paper.
@@ 6 Acknowledgments
S1 Constructed exceptional network
@@
S2 Redundant regulators
In this section, we address potential issues arising from discrepancies between the dynamics of a regulatory network and its graph representation. These issues stem from redundant expressions in the updating rules that can make the network present inoperative edges with no causal effect.
Previous work has shown that inoperative edges are rare in the biological networks, but not so in random networks [48]. We have examined the most apparent inoperative connections in the Cell Collective database and in the ensembles of random networks we generated. Additional networks from Ref. [25] have been curated similarly. However, we cannot guarantee that all inoperative edges have been eliminated. We aim to explain why, for the purposes of our study, a more meticulously curated database is not necessary.
Our primary objective is to elucidate why some regulatory networks are unexpectedly difficult to control. In this context, “difficult” refers to a significant deviation from the $\langle|\mathrm{CK}|\rangle\sim\log_{2}(r)$ scaling law in Eq. 1.
These issues can appear in several forms: inoperative edges, inoperative input nodes, and disconnected graphs. The latter two issues are more visible manifestations of inoperative edges. In the following, we will demonstrate that, in all three cases, these undesirable features do not make the network more difficult to control than anticipated. It would be possible but nontrivial to remove all inoperative edges before the analysis, so we choose not to do so here because it will not affect our conclusions.
Inoperative edges: For each network, we evaluate $\langle|\mathrm{CK}|\rangle$ and $r$ by explicitly solving the dynamics. Our analysis therefore does not depend directly on network topology except for our removal of peripheral trees to find the network core and compute $n_{c}$.
Inoperative edges could result in an overestimate of $n_{c}$, due to inoperative edges that hide output peripheral trees by making them look like they influence the core. For a given network, if we compute a number of isolated fixed points $s>0$ while including possible inoperative edges, then we know that our measurement of $n_{c}$ is correct: note that extra peripheral nodes created by inoperative edges would lead to basins of attraction with size $>1$ and therefore would mean that we would measure $s=0$. If we fixed this issue by testing for and removing inoperative edges, this would only affect estimated control kernel sizes for cases in which we computed $s=0$ (since the extra edges would have the effect of making it appear that $s=0$ in cases where $s$ is actually greater than 0.) We find empirically that difficult control occurs only in networks for which we computed $s>0$, so removing inoperative edges would not affect our conclusions about difficult control.
Inoperative input nodes: In cases where a node has neither real input nor output edges, it acts as an inoperative input node. This results in an artificial duplication of the number of attractors, as the inoperative input node introduces two possible entries (0/1) that are not present in the states of a more accurately curated network. Our algorithm for determining $\langle|\mathrm{CK}|\rangle$ will automatically include this irrelevant input, thus increasing $\langle|\mathrm{CK}|\rangle$ by 1, consistent with the $\langle|\mathrm{CK}|\rangle\sim\log_{2}(r)$ scaling law.
Disconnected graphs: Inoperative edges might obscure the fact that, from a dynamical perspective, the network can be reduced to two independent dynamical systems. After removing all inoperative edges, the network representation would correspond to two or more disconnected graphs. In this case, the control kernel for each global attractor must include all controlling nodes from the control kernels of the individual, fully connected components. The total number of attractors for the entire system would be the product of the attractors from each component. The $\langle|\mathrm{CK}|\rangle\sim\log_{2}(r)$ scaling law remains valid due to the product rule of logarithms. (The inoperative input nodes in the previous scenario are the simplest example of disconnected subgraphs consisting of a single node.)
S3 Exploring the shapes of residual distributions
In this section, we examine the statistical distributions of our predictions. As in Figure 5B and C, we consider the distribution of the residual differences between the actual $\langle|\textrm{CK}|\rangle$ and our predictions using both our simple logarithmic scaling (Eq. 1) and our corrected scaling (Eq. 2).
Figures 5B and C provide a visual confirmation that our corrected expressions for $\langle|\textrm{CK}|\rangle$ represent an improvement over the simpler estimates. Specifically, the distribution of the corrected residuals shows greater consistency with zero, with a smaller variance, due to a reduction in the predictions asymmetrically contributing to the right tail of our distribution. Here, we provide an estimate of the improvement this feature offers.
Before adopting any test for deviation from normality, it is important to emphasize that we have no reason to expect the distribution of our predicted residuals to follow a normal distribution, whether using our simpler or corrected model. We can still use such tests to quantify the reduction in the weight of the residual tails relative to any Gaussian fit. If our corrected approximation reduces the number of data points in the right tail of the distribution, this would manifest as a significant change in the test statistic of a normality test, e.g. a Shapiro-Wilk test as in the following. We indeed observe this effect. We analyzed the four distributions of residuals that correspond to the choice of biological vs. random networks together with the choice of our simpler vs. corrected scaling. The results of our test are as follows: the Shapiro-Wilk test statistic ($W$) increased from 0.64 to 0.85 for the biological network, and from 0.68 to 0.87 for the random networks. In both cases the test statistic increases by about $30\%$. (We also observe a corresponding increase in the p-values, which we should disregard, as our goal is not to validate the null hypothesis of normality.)
A more meaningful interpretation comes from complementing these changes in the test statistic with a visual inspection of the quantiles of our distributions vs. the quantiles of a normal distribution. These Q–Q plots, for all four cases discussed here, can be found in Figure S2, where the vertical axes (ordered quantities) represent our predictions, and the horizontal axes (theoretical quantities) show the expectation that would correspond to normally distributed residuals. Perfect normality of our distributions would correspond to a linear scaling of our data with respect to the theoretical quantiles, aligning with the red line shown in each panel.
For both biological and random networks, the dominant source of deviation from normality is represented by the largest ordered values in our distribution. This is the same result already shown in Figure 5, magnified here by revisiting it in terms of how much departure from a normal distribution they would contribute to if a normal distribution were the expectation.
One notable observation from the Q–Q plots is that the remaining primary source of non-normality in our distributions is the higher frequency with which our predictions are more accurate than a normal distribution would suggest once fitted to accommodate the tails of our distribution of residuals. This is evident in our figures as the horizontal plateau at ordered values equal to zero, where the number of cases with $\langle|\textrm{CK}|\rangle$ approximately zero occurs much more often than expected by a normal distribution. Our corrected approximation preserves this positive feature and decreases the discrepancies at larger observed values by reducing the instances where the simpler scaling underestimates the actual $\langle|\textrm{CK}|\rangle$.
S4 Additional supplemental figures
Figure S1: Mean control kernel sizes as a function of the logarithm of the number of attractors for three random network ensembles. Here we replot Fig. 4B to emphasize differences that arise due to constructing random networks in different ways: using random truth tables (dark triangles), threshold networks with balanced threshold (light squares), and threshold networks with zero threshold (green circles). Note that all the outliers identified as triangles in Fig. 4B come from the zero threshold ensemble. This figure omits 4 networks that were included in Fig. 9A of Ref. [21] due to our more strict threshold that excludes networks that have uncontrollable cycles with large basins.
Figure S2: Quantile–quantile (Q–Q) plots of the residual distributions shown in Fig. 5B and 5C. In all four cases displayed, the vertical axes (ordered quantiles) represent the observed residuals, and the horizontal axes (theoretical quantiles) represent the expected values assuming normally distributed residuals. A perfect match to normality would align with the red diagonal line, indicating linear scaling. The two plots on the left correspond to predictions using logarithmic scaling, while the plots on the right use our corrected scaling. The top row shows results for biological networks, and the bottom row shows results for random networks.
@@
name size edges core nodes mean CK size attr. isol. f.p.
$n$ $|E|$ $n_{\mathrm{c}}$ $|CK|\right>$ $r$ $s$
Arabidopsis Thaliana Cell Cycl… 14 66 14 0 1 0
B Cell Differentiation 22 44 20 5.9 61 0
BT474 Breast Cell Line Long-te… 25 74 25 14 3344 256
BT474 Breast Cell Line Short-t… 16 51 16 9.5 253 24
Body Segmentation In Drosophil… 17 31 12 3.4 10 0
Budding Yeast Cell Cycle 20 46 20 4.8 33 0
Budding Yeast Cell Cycle 2009 18 59 18 0 1 0
CD4+ T Cell Differentiation An… 18 84 18 8.6 324 0
Cardiac Development 15 39 15 2.8 6 0
Cell Cycle Transcription By Co… 9 19 9 1 2 0
Cortical Area Development 5 14 5 1 2 0
Death Receptor Signaling 28 48 25 4.7 97 0
FGF Pathway Of Drosophila Sign… 23 33 9 9 512 512
Fanconi Anemia And Checkpoint … 15 66 15 0 1 0
Guard Cell Abscisic Acid Signa… 44 82 37 5.4 95 0
HCC1954 Breast Cell Line Long-… 25 74 25 17.8 9452 1504
HCC1954 Breast Cell Line Short… 16 51 16 9 274 10
HH Pathway Of Drosophila Signa… 24 45 13 13 8192 8192
Human Gonadal Sex Determinatio… 19 79 18 2.3 3 0
Iron Acquisition And Oxidative… 22 40 17 2 4 0
Lac Operon 13 25 13 3.2 9 0
Mammalian Cell Cycle 20 52 18 1.7 3 0
Mammalian Cell Cycle 2006 10 35 10 0 1 0
Metabolic Interactions In The … 12 31 11 6.2 40 0
Neurotransmitter Signaling Pat… 16 22 14 2 4 0
Oxidative Stress Pathway 19 34 17 1 2 0
Predicting Variabilities In Ca… 15 39 15 2.8 6 0
Pro-inflammatory Tumor Microen… 26 81 26 2.5 6 0
Processing Of Spz Network From… 24 34 6 6 64 64
Regulation Of The L-arabinose … 13 21 10 4.1 21 0
SKBR3 Breast Cell Line Long-te… 25 85 25 17.3 3185 960
SKBR3 Breast Cell Line Short-t… 16 46 16 10.7 702 104
Septation Initiation Network 31 57 29 9.4 2240 0
T Cell Differentiation 23 38 23 5 33 0
T-Cell Signaling 2006 40 55 10 3.1 10 0
T-LGL Survival Network 2011 Re… 18 43 17 2.3 3 0
TOL Regulatory Network 24 58 10 10 1024 1024
Toll Pathway Of Drosophila Sig… 11 13 2 2 4 4
Treatment Of Castration-Resist… 42 65 14 14 16384 16384
Trichostrongylus Retortaeformi… 26 59 25 3.2 12 0
Tumour Cell Invasion And Migra… 32 158 22 3.3 15 0
Vegf Pathway Of Drosophila Sig… 18 26 8 8 256 256
Wg Pathway Of Drosophila Signa… 26 43 23 14 16384 0
Yeast Apoptosis 73 126 30 13.3 11520 0
@@
PubMed ID size edges core nodes mean CK size attr. isol. f.p.
$n$ $|E|$ $n_{\mathrm{c}}$ $|CK|\right>$ $r$ $s$
1753781 8 30 8 3.363636364 11 1
11082279 11 20 11 2 4 0
19524598 12 26 12 2 4 0
20169167 7 12 7 1 2 0
20659480 20 30 13 1.666666667 4 0
21563979 13 25 13 3.222222222 9 0
21639591 25 57 25 2 4 0
21853041 11 30 11 2.666666667 8 0
22192526 9 19 8 3 7 0
23056457 17 41 16 3.833333333 12 0
23469179 10 23 9 3.125 8 0
23764028 5 6 2 2 4 4
24376455 10 24 9 2 4 0
25163068 30 44 27 7.265306122 241 0
25967891 13 42 12 4.1 10 0
26244885 30 54 29 8.397435897 1104 0
26346668 30 42 2 2 4 4
27613445 25 68 23 3.6 22 0
28187161 30 54 27 6.823529412 220 0
28209158 9 34 9 1.666666667 3 0
28455685 15 24 15 2 4 0
29186334 29 83 16 5.238095238 25 0
29596489 23 55 18 2.25 5 0
30024932 21 62 16 2.333333333 6 0
30104572 5 12 5 2 3 0
30148917 18 30 14 4 16 0
30323768 21 36 16 3.8 18 0
30530226 23 39 19 2 5 0
30546316 30 58 21 10.60769231 1056 0
31949240 24 46 21 5.43902439 45 0
32054948 27 80 24 2.25 6 0
19622164_TGF_alpha 17 68 9 2.8 7 0
19622164_TGF_beta1 17 68 9 2.8 7 0
20221256_manual_mod 11 27 11 3.615384615 22 0
21450717_model_2_brassinosteroid 10 12 3 3 8 8
21450717_model_4_2_BRX_integration 20 30 17 3 14 0
21450717_model_5_1_BRX_integration 20 31 17 2 4 0
21450717_model_5_2 20 31 17 2 4 0
23169817_high_dna_damage 16 60 16 0 1 0
23169817_low_dna_damage 16 60 16 0 1 0
23630177_metastatic_melanoma 10 25 9 2 4 0
23658556_model_1 14 35 14 3.777777778 10 0
23658556_model_2 16 38 16 3.2 9 0
23658556_model_3 15 38 15 3.75 9 0
23658556_model_4 17 42 17 3 10 0
23658556_model_5 15 37 15 3.125 8 0
23658556_model_6 15 37 15 3.125 8 0
23658556_model_7 17 40 17 4.5 14 0
23658556_model_8 16 38 16 4.1 12 0
23658556_model_9 17 42 17 3.25 9 0
24564942_threshold 8 23 8 2.6 5 0
28426669_ARF10_greater_ARF5 16 39 15 3.3 14 0
28426669_ARF10_smaller_ARF5 16 39 15 3.3 14 0